OpenStudy (anonymous):

find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch side

OpenStudy (haltiamali):

If it's equilateral than all the measurements are the same.(: Do you know the area of a triangle?

OpenStudy (anonymous):

\[A=\frac{ \sqrt{3} }{ 4 }\times side ^{2}\]

OpenStudy (anonymous):

\[\frac{ \sqrt{3} }{ 4 } \times 36\]\[.433\times36=15.58\]

OpenStudy (anonymous):

The equation is Area= 1/2 (anthom)(perimeter)